OpenStudy (anonymous):

The temperature was 5 degrees F in the afternoon and dropped 10 degrees F. What is the expression needed to determine the new temperature? What is the new temperature?

OpenStudy (mathmale):

Should you be using addition, subtraction, multiplication or division?

OpenStudy (anonymous):

idk

OpenStudy (xhawaiiangirlx):

At first the temperature was 5 degrees right? Then it dropped to 10 degrees so when you're trying to drop a number to get to the next what do you do?

OpenStudy (xhawaiiangirlx):

@XbubblegummieX

OpenStudy (anonymous):

Would you subtract it ?

OpenStudy (xhawaiiangirlx):

correct!

OpenStudy (xhawaiiangirlx):

So if you want to make an expression with subtraction what would it be?

OpenStudy (xhawaiiangirlx):

Since it started at 5 degrees and it dropped 10 what would your expression be? Hint: the answer becomes a negative.

OpenStudy (anonymous):

5 - (-10) ?

OpenStudy (mathmale):

Why the double " - " ? Please note that 5-(-10)=5+10=15.

OpenStudy (anonymous):

that's the negative for the 10